140827 LML Observations

Isaac arrived at the collector at 9:30am. The doors to the collector had been deployed however, no sample had been collected. The collector was then disassembled for cleaning. This was preformed by utilizing 409 and DI water. Upon completion of this process the collector was then reassembled and reset for the next fog event.

Leave a Reply

Your email address will not be published. Required fields are marked *